# Limits of Absoluteness of Observed Events in Timelike Scenarios: A No-Go Theorem

### v1

@@ IV.3 Impossibility of discarding the Absoluteness of Pseudo Events
Above, we showed that we can split the AOE assumption into EOM and EJPD, and still obtain a contradiction with quantum mechanics if we replace AOE with EOM, so long as we strengthen SPE into OM. We now argue that we can split EOM up further, into Absoluteness of Truly-Observed Events ATOE, and Absoluteness of Pseudo Events APE defined respectively as,
Definition 3 (Absoluteness of Truly-Observed Events (ATOE)).
@@
Definition 4 (Absoluteness of Pseudo Events (APE)).
@@
We can summarise the connections between the assumptions discussed in Sec. IV.3 and Sec. IV.2 as follows:
$\begin{split}\textbf{EOM}&+\textbf{OM}\implies\textbf{OPEM}\\ \textbf{EOM}+\textbf{NRC}&\implies\textbf{ATOE}+\textbf{APE}\end{split}$ (48)
The second implication in Eq. (48) is due to the fact that the assumption EOM already assumes that $p(c,d|x,y)$ exists ( i.e.,APE). Moreover, the second part of EOM is that $p(a|x,c)$ and $p(b|y,c,d)$ exists. Then together this implies that there exists a probability $p(a,b|x,y)$ as
@@
$=\sum_{c,d|x,y}p(c,d)p(a|x,c,d)p(b|y,c,d)$
which is what ATOE is. The second line of the above equation is due to NRC.
Having shown in Section IV.2 that we can remove the EJPD assumption while still being able to generate an inequality which quantum mechanics can violate, we now wish to consider whether we can further remove assumption, ideally to see if ATOE alone is enough to produce an inequality which Quantum Mechanics violates. This question is motivated by prima facie APE being violated by standard quantum mechanics, given quantum mechanics describes Charlie’s (Debbie’s) lab as being in a superposition from Alice’s (Bob’s) perspective until they open the lab, and given Alice (Bob) can unitarily undo Charlie’s (Debbie’s) measurement - which requires the measurement itself to have been unitary according to quantum mechanics in the first place.
To answer this question, let us consider the case without APE, where we have the following factorised form of observed probabilities,
@@
We can no longer factor this as a single convex average $\sum_{c,d}p(c,d)\,S_{c,d}$ as done while establishing inequality Eq. (35). The pointwise bound $|S^{c,d}|\leq 2$ for fixed $c,d$ does not, by itself, constrain mixtures with different context-dependent weights. As a result, the overall value of $S$ derived without assuming APE may even exceed the quantum bound, undermining the no-go theorem.
@@
which implies S=1+1+1(1)=4S=1+1+1-(-1)=4, violating not only the bound provided in Eq. (35) but also the quantum bound $S_{Q}=2\sqrt{2}$, reaching the Boxworld bound $S_{B}=4$ [18]. This shows that without mediator independence (derived by combining ATS+NRC (Lemma 1) with AOE), it is not possible to establish the no-go result.
@@ V.1 Comparing to Related Results
In a prepare-and-measure scenario the connection between preparation contextuality and AOE is formulated through what they refer to as the Operational Friendliness no-go theorem. Crucially, Leifer and Pusey proved that, in a timelike prepare–and–measure scenario, imposing Time Symmetry together with No Retrocausality and $\lambda$-mediation yields the mediator/epistemic independence condition (e.g., $p(\lambda|x)=p(\lambda)$ for the no-signalling preparations) usually assumed in preparation-noncontextuality arguments [17]. Consequently these assumptions together imply a timelike Bell-CHSH constraint which quantum theory violates. In contrast, the Operational Friendliness no-go theorem derives this contradiction with quantum theory by combining AOE with Operational Agency- a stronger version of operational equivalence used in the discussion of Spekkens’ contextuality[38, 51, 52]. By using similar operational correlations involving time-ordered measurements, we employ AOE to establish our Causal Friendliness no-go result through a different yet equally powerful route. We therefore view our Causal Friendliness no-go theorem as complementary to the Operational Friendliness no-go theorem. However, the two are not hierarchically ordered because Time Symmetry and preparation-noncontextuality are logically different principles and can be independent in other scenarios.
Note that the Causal Friendliness constraints are less strict than the generalised-noncontextuality constraints for the sequential measurement scenario. However, for the bipartite scenario we discuss here, where each party can choose between two dichotomic measurements (the $2222$-Bell scenario), noncontextuality constraints provide the same Bell-CHSH-like inequality. To prove that our constraint is strictly weaker than those noncontextual constraints, one would need to consider scenarios beyond the $2222$-Bell scenario, where one can violate a Causal Friendliness inequality but not an inequality derived constraints similar to those used in proofs of Kochen-Specker contextuality [52] or preparation contextuality [51]. We leave the derivation of such Causal Friendliness inequalities beyond the $2222$-Bell scenario for future work.
@@ V.2 Implications of the findings for interpretational issues
In the dBB framework, every system is associated with a set of ontic variables (the system’s exact position) that evolves under a guiding equation. These variables assign definite outcomes to all measurement events, so both the Truly-Observed Events and the Pseudo Events possess ontic definiteness in every run. In the terminology of this work, dBB explicitly maintains ATOE and APE. However, as the quantum potential is explicitly nonlocal (allowing measurement results to depend on factors outside the past light-cone of the measurement), in the Time Symmetric sector, dBB circumvents our no-go results by abandoning the condition of SPE.
On the other hand, the Many-Worlds Interpretation (MWI) fundamentally rejects the very notion of single, absolute outcomes. In the MWI, a measurement does not yield a unique, classical result. Rather, it induces a branching of the global wavefunction into dynamically independent components, each realising a distinct record of the possible outcomes. Consequently, the MWI explicitly denies the AOE assumption on which our no-go theorem is based. From the perspective of an MWI adherent, the apparent incompatibility between quantum predictions and the Causal-Friendliness inequality simply reflects the fact that the observed events in any given branch represent only one relative component of a larger, branch-dependent reality. In this view, the violation of AOE at the operational level arises naturally from the multiplicity of branching records in the universal wavefunction, thereby making the Causal-Friendliness no-go result completely compatible with the logic used in MWI.
@@ V.3 Summary and future directions
In this paper, building on recent results, we introduced the Causal Friendliness Paradox, a time-ordered analogue of Bong et al ’s Local Friendliness no-go theorem. To generate this no-go theorem, we replace the usual locality assumption with Axiological Time Symmetry (ATS), showing that, when combined with Absoluteness of Observed Events (AOE), No Retrocausality (NRC), and Screening via Pseudo Events (SPE), we obtain a causal inequality. We then showed that quantum mechanics violates this inequality, demonstrating quantum statistics must be incompatible with at least one of these assumptions. To probe which assumption might be incompatible, we then examine whether AOE in its entirety is essential for this no-go result. We propose a weaker, operational form of AOE that still leads to the inequality that quantum mechanics violates. This result shows that even under relaxed assumptions, quantum theory resists reconciliation with classical notions of absolute events, reinforcing the foundational significance of Wigner’s Friend-type paradoxes in timelike scenarios. However, we finally showed that trying to remove the absoluteness of the individual Pseudo Events $c$ and $d$ undermines our ability to arrive at a contradiction with quantum mechanics, allowing us to not only surpass the maximum quantum value, but reach the maximum possible algebraic value for the correlators for the scenario. This illustrates the crucial role played by the Absoluteness of Pseudo Events in establishing the no-go result.
Our findings opens several directions for the future work, such as extending the weakening the Absoluteness of Observed Events to derive other no-go theorems based on different set of assumptions (e.g. macrorealism [57]), as well as exploring experimentally accessible scenarios that could test genuine Causal-Friendliness violations. A further direction is to relate the present framework to studies of the quantum-to-classical transition. Since the framework of weakening of absoluteness of events specifies when events acquire effective classicality, it may offer an operational perspective on when and how definite outcomes emerge through decoherence [58, 59] and coarse-grained measurements [60].

### v2

@@ IV.3 Impossibility of discarding the Absoluteness of Pseudo Events
Above, we showed that we can weaken the Absoluteness of Observed Events assumption to Existence of marginals, and still obtain a contradiction with quantum mechanics, so long as we strengthen Screening via Pseudo Events into Operational mediation. We now argue that we can split Existence of marginals up further, into Absoluteness of Truly-Observed Events , and Absoluteness of Pseudo Events defined respectively as,
Definition 6 (Absoluteness of Truly-Observed Events).
@@
Definition 7 (Absoluteness of Pseudo Events).
@@
We can summarise the connection between the assumptions discussed in Sec. IV.3 and Sec. IV.2 as follows:
$\textbf{Existence of marginals}+\textbf{No Retrocausality}$
$\implies\textbf{Absoluteness of Truly-Observed Events}$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\textbf{Absoluteness of Pseudo Events}$ (52)
The implication in Eq. (IV.3) is due to the fact that the assumption Existence of marginals already assumes that $p(c,d|x,y)$ exists ( i.e.,Absoluteness of Pseudo Events). Moreover, the second part of Existence of marginals is that $p(a|x,c)$ and $p(b|y,c,d)$ exists. Then together this implies that there exists a probability $p(a,b|x,y)$ as
@@
$=\sum_{c,d}p(c,d)p(a|x,c,d)p(b|y,c,d)$
which is what Absoluteness of Truly-Observed Events is. The second line of the above equation is due to No Retrocausality. We do not claim that this particular implication will hold generally, but arises from the particular way events, Pseudo Events, and operational accessibility are defined in this Causal Time-Symmetric Friendliness scenario. This is because No Retrocausality is a causal assumption, whereas Absoluteness of Truly-Observed Events and Absoluteness of Pseudo Events are not themselves causal statements, and no general implication between them is expected outside the present framework.
Having shown in Section IV.2 that we can remove the Existence of Joint Probability assumption while still being able to generate an inequality which quantum mechanics can violate, we now wish to consider whether we can further remove assumption, ideally to see if Absoluteness of Truly-Observed Events alone is enough to produce an inequality which Quantum Mechanics violates. This question is motivated by prima facie Absoluteness of Pseudo Events being violated by standard quantum mechanics, given quantum mechanics describes Charlie’s (Debbie’s) lab as being in a superposition from Alice’s (Bob’s) perspective until they open the lab, and given Alice (Bob) can unitarily undo Charlie’s (Debbie’s) measurement - which requires the measurement itself to have been unitary according to quantum mechanics in the first place. By the term “standard quantum mechanics”, here we mean the description of quantum mechanics that follows from, or which is mathematically equivalent to, the Copenhagen interpretation [5].
To answer this question, let us consider the case without Absoluteness of Pseudo Events, where we have the following factorised form of observed probabilities,
@@
We can no longer factor this as a single convex average $\sum_{c,d}p(c,d)\,S_{c,d}$ as done while establishing inequality Eq. (37). The pointwise bound $|S^{c,d}|\leq 2$ for fixed $c,d$ does not, by itself, constrain mixtures with different context-dependent weights. As a result, the overall value of $S$ derived without assuming Absoluteness of Pseudo Events may even exceed the quantum bound, undermining the no-go theorem.
@@
which implies S=1+1+1(1)=4S=1+1+1-(-1)=4, violating not only the bound provided in Eq. (37) but also the quantum bound $S_{Q}=2\sqrt{2}$, reaching the Boxworld bound $S_{B}=4$ [47]. This shows that without mediator independence (derived by combining Axiological Time Symmetry+No Retrocausality (Lemma 1) with Absoluteness of Observed Events), it is not possible to establish the no-go result.
@@ V.1 Comparing to Related Results
In a prepare-and-measure scenario the connection between preparation contextuality and Absoluteness of Observed Events is formulated through what they refer to as the Operational Friendliness no-go theorem. Crucially, Leifer and Pusey proved that, in a timelike prepare–and–measure scenario, imposing Time Symmetry together with No Retrocausality and $\lambda$-mediation yields the mediator/epistemic independence condition (e.g., $p(\lambda|x)=p(\lambda)$ for the no-signalling preparations) usually assumed in preparation-noncontextuality arguments [39]. Consequently these assumptions together imply a timelike Bell-CHSH constraint which quantum theory violates. In contrast, the Operational Friendliness no-go theorem derives this contradiction with quantum theory by combining Absoluteness of Observed Events with Operational Agency- a stronger version of operational equivalence used in the discussion of Spekkens’ contextuality[56, 61, 62]. By using similar operational correlations involving time-ordered measurements, we employ Absoluteness of Observed Events to establish our Causal Time-Symmetric Friendliness no-go result through a different yet equally powerful route. We therefore view our Causal Time-Symmetric Friendliness no-go theorem as complementary to the Operational Friendliness no-go theorem. However, the two are not hierarchically ordered because Time Symmetry and preparation-noncontextuality are logically different principles and can be independent in other scenarios.
Note that it is worth investigating the difference between the Causal Time-Symmetric Friendliness constraints with the generalised-noncontextuality constraints for the sequential measurement scenario. However, for the bipartite scenario we discuss here, where each party can choose between two dichotomic measurements (the $2222$-Bell scenario), noncontextuality constraints provide the same Bell-CHSH-like inequality. To determine whether our constraints are weaker or stronger than those imposed by noncontextuality, one would need to consider scenarios beyond the $2222$-Bell scenario, in which a Causal Time-Symmetric Friendliness inequality can be violated while inequalities derived from Kochen–Specker contextuality [62] or preparation contextuality [61] cannot, and vice versa. We leave the derivation of such Causal Time-Symmetric Friendliness inequalities beyond the $2222$-Bell scenario for future work.
@@ V.2 Implications of the findings for interpretational issues
In the dBB framework, every system is associated with a set of ontic variables (the system’s exact position) that evolves under a guiding equation. These variables assign definite outcomes to all measurement events, so both the Truly-Observed Events and the Pseudo Events possess ontic definiteness in every run. In the terminology of this work, dBB explicitly maintains Absoluteness of Truly-Observed Events and Absoluteness of Pseudo Events. However, as the quantum potential is explicitly nonlocal (allowing measurement results to depend on factors outside the past light-cone of the measurement), in the Time Symmetric sector  [^2] In general, dBB theory, like standard quantum theory, does not satisfy Operational Time Symmetry. However, one may consistently restrict attention to a time-symmetric sector of the theory [59, 53]., dBB circumvents our no-go results by abandoning the condition of Screening via Pseudo Events.
On the other hand, the Many-Worlds Interpretation (MWI) fundamentally rejects the very notion of single, absolute outcomes. In the MWI, a measurement does not yield a unique, classical result. Rather, it induces a branching of the global wavefunction into dynamically independent components, each realising a distinct record of the possible outcomes. Consequently, the MWI explicitly denies the Absoluteness of Observed Events assumption on which our no-go theorem is based. From the perspective of an MWI adherent, the apparent incompatibility between quantum predictions and the Causal Time-Symmetric Friendliness inequality simply reflects the fact that the observed events in any given branch represent only one relative component of a larger, branch-dependent reality. In this view, the violation of AOE at the operational level arises naturally from the multiplicity of branching records in the universal wavefunction, thereby making the Causal Time-Symmetric Friendliness no-go result completely compatible with the logic used in MWI.
@@ V.3 Summary and future directions
In this paper, building on recent results, we introduced the Causal Time-Symmetric Friendliness Paradox, a time-ordered analogue of Bong et al ’s Local Friendliness no-go theorem. To generate this no-go theorem, we replace the usual locality assumption with Axiological Time Symmetry (ATS), showing that, when combined with Absoluteness of Observed Events (AOE), No Retrocausality (NRC), and Screening via Pseudo Events (SPE), we obtain a causal Time-Symmetric Friendliness inequality. We then showed that quantum mechanics violates this inequality, demonstrating quantum statistics must be incompatible with at least one of these assumptions. To probe which assumption might be incompatible, we then examine whether AOE in its entirety is essential for this no-go result. We propose a weaker, operational form of AOE that still leads to the inequality that quantum mechanics violates. This result shows that even under relaxed assumptions, quantum theory resists reconciliation with classical notions of absolute events, reinforcing the foundational significance of Wigner’s Friend-type paradoxes in timelike scenarios. However, we finally showed that trying to remove the absoluteness of the individual Pseudo Events $c$ and $d$ undermines our ability to arrive at a contradiction with quantum mechanics, allowing us to not only surpass the maximum quantum value, but reach the maximum possible algebraic value for the correlators for the scenario. This illustrates the crucial role played by the Absoluteness of Pseudo Events in establishing the no-go result.
Our findings opens several directions for the future work, such as extending the weakening the Absoluteness of Observed Events to derive other no-go theorems based on different set of assumptions (e.g. macrorealism [36]), as well as exploring experimentally accessible scenarios that could test genuine Causal Time-Symmetric Friendliness violations. A further direction is to relate the present framework to studies of the quantum-to-classical transition. Since the framework of weakening of absoluteness of events specifies when events acquire effective classicality, it may offer an operational perspective on when and how definite outcomes emerge through decoherence [70, 51] and coarse-grained measurements [34].